DVIPS: A TEX Driver

                   Tomas Rokicki, rokicki@cs.stanford.edu

FAX: (415) 327–3329 [This manual describes version 5.58.]

The dvips program converts a TEX dvi file into a PostScript file for printing or distribution. Seldom has such a seemingly easy programming task required so much effort.

1. Why Use dvips?

The dvips program has a number of features that set it apart from other PostScript drivers for TEX. This rather long section describes the advantages of using dvips, and may be skipped if you are just interested in learning how to use the program. Installation is covered in section 14 near the end of this document.

The dvips driver generates excellent, standard PostScript, that can be included in other documents as figures or printed through a variety of spoolers. The generated PostScript requires very little printer memory, so very complex documents with a lot of fonts can easily be printed even on PostScript printers without much memory, such as the original Apple LaserWriter. The PostScript output is also compact, requiring less disk space to store and making it feasible as a transfer format.

Even those documents that are too complex to print in their entirety on a particular printer can be printed, since dvips will automatically split such documents into pieces, reclaiming the printer memory between each piece.

The dvips program supports graphics in a natural way, allowing PostScript graphics to be included and automatically scaled and positioned in a variety of ways.

Printers with resolutions other than 300 dpi are also supported, even if they have a different resolution in the horizontal and vertical directions. High resolution output is supported for typesetters, including an option that compresses the bitmapped fonts so that typesetter virtual memory is not exhausted. This option also significantly reduces the size of the PostScript file and decoding in the printer is very fast.

Missing fonts can be automatically generated if exists on the system, or fonts can be converted from gf to pk format on demand. If a font cannot be generated, a scaled version of the same font at a different size can be used instead, although dvips will complain loudly about the poor æsthetics of the resulting output.

DVIPS: A TEX Driver 2

Users will appreciate features such as collated copies and support for tpic, psfig, emtex, and METAPOST; system administrators will love the support for multiple printers, each with their own configuration file, and the ability to pipe the output directly to a program such as lpr. Support for MS-DOS, OS/2, and VMS in addition to UNIX is provided in the standard distribution, and porting to other systems is easy.

One of the most important features is the support of virtual fonts, which add an entirely new level of flexibility to TEX. Virtual fonts are used to give dvips its excellent PostScript font support, handling all the font remapping in a natural, portable, elegant, and extensible way. The dvips driver even comes with its own afm2tfm program that creates the necessary virtual fonts and TEX font metric files automatically from the Adobe font metric files.

Source is provided and freely distributable, so adding a site-specific feature is possible. Adding such features is made easier by the highly modular structure of the program.

There is really no reason to use another driver, and the more people use dvips, the less time will be spent fighting with PostScript and the more time will be available to create beautiful documents. So if you don’t use dvips on your system, get it today.

2. Using dvips

To use dvips, simply type

    localhost> dvips foo

where foo.dvi is the output of TEX that you want to print. If dvips has been installed correctly, the document should come out of your default printer.

If you use fonts that have not been used on your system before, they may be automatically generated; this process can take a few minutes. The next time that document is printed, these fonts will already exist, so printing will go much faster.

Many options are available; they are described in a later section. For a brief summary of available options, just type

    localhost> dvips

3. Paper Size and Landscape Mode

Most TEX documents at a particular site are designed to use the standard paper size (for example, letter size in the United States or A4 in Europe.) The dvips program defaults to these paper sizes and can be customized for the defaults at each site or on each printer.

But many documents are designed for other paper sizes. For instance, you may want to design a document that has the long edge of the paper horizontal. This can be useful when typesetting booklets, brochures, complex tables, or many other documents. This type of paper orientation is called landscape orientation (the ‘normal’ orientation is portrait). Alternatively, a document might be designed for ledger or A3 paper.

Since the intended paper size is a document design decision, and not a decision that is made at printing time, such information should be given in the TEX file and not on the dvips command line. For this reason, dvips supports a papersize special. It is hoped that this special will become standard over time for TEX previewers and other printer drivers.

The format of the papersize special is

    \special{papersize=8.5in,11in}

where the dimensions given above are for a standard letter sheet. The first dimension given is the horizontal size of the page, and the second is the vertical size. The dimensions supported are the same as for TEX; namely, in (inches), cm (centimeters), mm (millimeters), pt (points), sp (scaled points), bp (big points, the same as the default PostScript unit), pc (picas), dd (didot points), and cc (ciceros).

For a landscape document, the papersize comment would be given as

    \special{papersize=11in,8.5in}

An alternate specification of landscape is to have a special of the form

    \special{landscape}

This is supported for backward compatibility, but it is hoped that eventually the papersize comment will dominate.

Of course, using such a command only informs dvips of the desired paper size; you must still adjust the hsize and vsize in your TEX document to actually use the full page.

The papersize special must occur somewhere on the first page of the document.

The dvips default landscape configuration is presently as though the paper were rotated ninety degrees counterclockwise; this seems to be the convention adopted by previewers that the author is familiar with. If for some reason you need your landscape orientation to be rotated clockwise, simply include at the top of your TEX file or in some included macro file

    \special{!  /landplus90 true store}

to set this orientation. Alternatively, you can change the default value of landplus90 in the tex.lpro file before compilation, or include a header file that just includes the definition

/landplus90 true store.

4. Including PostScript Graphics

Scaling and including PostScript graphics is a breeze—if the PostScript file is correctly formed. Even if it is not, however, the file can usually be accommodated with just a little more work. The most important feature of a good PostScript file—from the standpoint of including it in another document—is an accurate bounding box comment.

4.1 The Bounding Box Comment

Every well-formed PostScript file has a comment describing where on the page the graphic is located, and how big that graphic is. This information is given in terms of the lower left and upper right corners of a box just enclosing the graphic, and is thus referred to as a bounding box. These coordinates are given in PostScript units (there are precisely 72 PostScript units to the inch) with respect to the lower left corner of the sheet of paper.

To see if a PostScript file has a bounding box comment, just look at the first few lines of the file. (PostScript is standard ASCII, so you can use any text editor to do this.) If within the first few dozen lines there is a line of the form

    %%BoundingBox: 0 1 2 3

(with any numbers), chances are very good that the file is Encapsulated PostScript and will work easily with dvips. If the file contains instead a line like

    %%BoundingBox: (atend)

the file is still probably Encapsulated PostScript, but the bounding box (that dvips needs to position the graphic) is at the end of the file and should be moved to the position of the line above. This can be done with that same text editor, or with a simple Perl script.

If the document lacks a bounding box altogether, one can easily be added. Simply print the file. Now, take a ruler, and make the following measurements. All measurements should be in PostScript units, so measure it in inches and multiply by 72. Alternatively, the bbfig program distributed with dvips in the contrib directory can be used to automate this process.

From the left edge of the paper to the leftmost mark on the paper is llx, the first number. From the bottom edge of the paper to the bottommost mark on the paper is lly, the second number. From the left edge of the paper to the rightmost mark on the paper is urx, the third number. The fourth and final number, ury, is the distance from the bottom of the page to the uppermost mark on the paper.

Now, add a comment of the following form as the second line of the document. (The first line should already be a line starting with the two characters ‘%!’; if it is not, the file probably isn’t PostScript.)

    %%BoundingBox: llx lly urx ury

Or, if you don’t want to modify the file, you can simply write these numbers down in a convenient place and use them when you import the graphic.

If the document does not have such a bounding box, or if the bounding box is given at the end of the document, please complain to the authors of the software package that generated the file; without such a line, including PostScript graphics can be tedious.

4.2 Using the EPSF Macros

Now you are ready to include the graphic into a TEX file. Simply add to the top of your TEX file a line like

    \input epsf

(or, if your document is in LaTEX or SliTEX, add the epsf style option, as was done to the following line).

    \documentstyle[12pt,epsf]{article}

This only needs to be done once, no matter how many figures you plan to include. Now, at the point you want to include the file, enter a line such as

    \epsffile{foo.ps}

If you are using LaTEX or SliTEX, you may need to add a \leavevmode command immediately before the \epsffile command to get certain environments to work correctly. If your file did not (or does not currently) have a bounding box comment, you should supply those numbers you wrote down as in the following example:

    \epsffile[100 100 500 500]{foo.ps}

(in the same order they would have been in a normal bounding box comment). Now, save your changes and run TEX and dvips; the output should have your graphic positioned at precisely the point you indicated, with the proper amount of space reserved.

The effect of the \epsffile macro is to typeset the figure as a TEX \vbox at the point of the page that the command is executed. By default, the graphic will have its ‘natural’ width (namely, the width of its bounding box). The TEX box will have depth zero and a ‘natural’ height. The graphic will be scaled by any dvi magnification in effect at the time.

Any PostScript graphics included by any method in this document (except bop-hook and its ilk) are scaled by the current dvi magnification. For graphics included with \epsffile where the size is given in TEX dimensions, this scaling will produce the correct, or expected, results. For compatibility with old PostScript drivers, it is possible to turn this scaling off with the following TEX command:

    \special{!  /magscale false def}

Use of this command is not recommended because it will make the \epsffile graphics the wrong size if global magnification is used in a dvi document, and it will cause any PostScript graphics to appear improperly scaled and out of position if a dvi to dvi program is used to scale or otherwise modify the document.

You can enlarge or reduce the figure by putting

    \epsfxsize=<dimen>

right before the call to \epsffile. Then the width of the TEX box will be <dimen> and its height will be scaled proportionately. Alternatively you can force the vertical size to a particular size with

    \epsfysize=<dimen>

in which case the height will be set and the width will be scaled proportionally. If you set both, the aspect ratio of the included graphic will be distorted but both size specifications will be honored.

By default, clipping is disabled for included EPSF images. This is because clipping to the bounding box dimensions often cuts off a small portion of the figure, due to slightly inaccurate bounding box arguments. The problem might be subtle; lines around the boundary of the image might be half their intended width, or the tops or bottoms of some text annotations might be sliced off. If you want to turn clipping on, just use the command

    \epsfclipon

and to turn clipping back off, use

    \epsfclipoff

A more general facility for sizing is available by defining the \epsfsize macro. You can redefine this macro to do almost anything. This TEX macro is passed two parameters by \epsffile. The first parameter is the natural horizontal size of the PostScript graphic, and the second parameter is the natural vertical size. This macro is responsible for returning the desired horizontal size of the graph (the same as assigning \epsfxsize above).

In the definitions given below, only the body is given; it should be inserted in

    \def\epsfsize#1#2{body}

Some common definitions are:

\epsfxsize This definition (the default) enables the default features listed above, by setting \epsfxsize to the same value it had before the macro was called.

0pt This definition forces natural sizes for all graphics by setting the width to zero, which turns on horizontal scaling.

#1 This forces natural sizes too, by returning the first parameter only (the natural width) and setting the width to it.

\hsize This forces all graphics to be scaled so they are as wide as the current horizontal size. (In LaTEX, use \textwidth instead of \hsize.)

0.5#1 This scales all figures to half of their natural size.

\ifdim#1>\hsize\hsize\else#1\fi This keeps graphics at their natural size, unless the width would be wider than the current \hsize, in which case the graphic is scaled down to \hsize.

If you want TEX to report the size of the figure as a message on your terminal when it processes each figure, give the command

    \epsfverbosetrue

4.3 Header Files

Often in order to get a particular graphic file to work, a certain header file might need to be sent first. Sometimes this is even desirable, since the size of the header macros can dominate the size of certain PostScript graphics files. The dvips program provides support for this with the header= special command. For instance, to ensure that foo.ps gets downloaded as part of the header material, the following command should be added to the TEX file:

    \special{header=foo.ps}

The dictionary stack will be at the userdict level when these header files are included.

For these and all other header files (including the headers required by dvips itself and any downloaded fonts), the printer VM budget is debited by some value. If the header file has, in its first 1024 bytes, a line of the form

    %%VMusage: min max

then the maximum value is used. If it doesn’t, then the total size of the header file in bytes is used as an approximation of the memory requirements.

4.4 Literal PostScript

For simple graphics, or just for experimentation, literal PostScript graphics can be included. Simply use a special command that starts with a double quote ("). For instance, the following (simple) graphic:

was created by typing:

    \vbox to 100bp{\vss % a bp is the same as a PostScript unit
    \special{" newpath 0 0 moveto 100 100 lineto 394 0 lineto
    closepath gsave 0.8 setgray fill grestore stroke}}

(Note that you are responsible for leaving space for such literal graphics.) Literal graphics are discouraged because of their nonportability.

4.5 Literal Headers

Similarly, you can define your own macros for use in such literal graphics through the use of literal macros. Literal macros are defined just like literal graphics, only you begin the special with an exclamation point instead of a double quote. These literal macros are included as part of the header material in a special dictionary called SDict. This dictionary is the first one on the PostScript dictionary stack when any PostScript graphic is included, whether by literal inclusion or through the \epsffile macros.

4.6 Other Graphics Support

There are other ways to include graphics with dvips. One is to use an existing package,

such as emtex, psfig, tpic, or METAPOST, all supported by dvips.

Other facilities are available for historical reasons, but their use is discouraged, in hope that some ‘sane’ form of PostScript inclusion shall become standard. Note that the main advantage of the \epsffile macros is that they can be adapted for whatever form of special eventually becomes standard, and thus only minor modifications to that one file need to be made, rather than revising an entire library of TEX documents.

Most of these specials use a flexible key and value scheme:

    \special{psfile=filename.ps[key=value]*}

This will download the PostScript file called filename.ps such that the current point will be the origin of the PostScript coordinate system. The optional key/value assignments allow you to specify transformations on the PostScript.

The possible keys are:

hoffset The horizontal offset (default 0)
voffset The vertical offset (default 0)
hsize  The horizontal clipping size (default 612)
vsize  The vertical clipping size (default 792)
hscale  The horizontal scaling factor (default 100)
vscale  The vertical scaling factor (default 100)
angle  The rotation (default 0)
clip  Enable clipping to the bounding box

The dimension parameters are all given in PostScript units. The hscale and vscale are given in non-dimensioned percentage units, and the rotation value is specified in degrees. Thus

    \special{psfile=foo.ps hoffset=72 hscale=90 vscale=90}

will shift the graphics produced by file foo.ps right by one inch and will draw it at 0.9 times normal size. Offsets are given relative to the point of the special command, and are unaffected by scaling or rotation. Rotation is counterclockwise about the origin. The order of operations is to rotate the figure, scale it, then offset it.

For compatibility with older PostScript drivers, it is possible to change the units that hscale and vscale are given in. This can be done by redefining @scaleunit in SDict by a TEX command such as

    \special{!  /@scaleunit 1 def}

The @scaleunit variable, which is by default 100, is what hscale and vscale are divided by to yield an absolute scale factor.

All of the methods for including graphics we have described so far enclose the graphic in a PostScript save/restore pair, guaranteeing that the figure will have no effect on the rest of the document. Another type of special command allows literal PostScript instructions to be inserted without enclosing them in this protective shield; users of this feature are supposed to understand what they are doing (and they shouldn’t change the PostScript graphics state unless they are willing to take the consequences). This command can take many forms, because it has had a tortuous history; any of the following will work:

    \special{ps:text}
    \special{ps::text}
    \special{ps::[begin]text}
    \special{ps::[end]text}

(with longer forms taking precedence over shorter forms, when they are used). Note that ps:: and ps::[end] do not do any positioning, so they can be used to continue PostScript literals started with ps: or ps::[begin]. There is also the command

    \special{ps: plotfile filename}

which will copy the commands from filename verbatim into the output (but omitting lines that begin with %). An example of the proper use of literal specials can be found in the file rotate.tex which makes it easy to typeset text turned 90 degrees.

To finish off this section, the following examples are presented without explanation:

\def\rotninety{\special{ps:currentpoint currentpoint translate 90
rotate neg exch neg exch translate}}\font\huge=cmbx10 at 14.4truept
\setbox0=\hbox to0pt{\huge A\hss}\vskip16truept\centerline{\copy0
\special{ps:gsave}\rotninety\copy0\rotninety\copy0\rotninety
\box0\special{ps:grestore}}\vskip16truept

A A A A

\vbox to 2truein{\special{ps:gsave 0.3 setgray}\hrule height 2in
width\hsize\vskip-2in\special{ps:grestore}\font\big=cminch\big
\vss\special{ps:gsave 1 setgray}\vbox to 0pt{\vskip2pt
\line{\hss\hskip4pt NEAT\hss}\vss}\special{ps:0 setgray}%
\hbox{\raise2pt\line{\hss NEAT\hss}\special{ps:grestore}}\vss}

Some caveats are in order when using the above forms. Make sure that each gsave on a page is matched with a grestore on the same page. Do not use save or restore. Use of these macros can interact with the PostScript generated by dvips if care is not taken; try to understand what the above macros are doing before writing your own. The \rotninety macro especially has a useful trick that appears again and again.

4.7 Dynamic Creation of PostScript Graphics Files

PostScript is an excellent page description language—but it does tend to be rather verbose. Compressing PostScript graphics files can often reduce them by more than a factor of five. For this reason, if the filename parameter to one of the graphics inclusion techniques starts with a backtick (`), the filename is instead interpreted as a command to execute that will send the actual file to standard output. Thus,

    \special{psfile="`zcat foo.ps.Z"}

will include the uncompressed version of foo.ps. Since such a command is not accessible to TEX, if you use this facility with the EPSF macros, you need to supply the bounding box parameter yourself, as in

    \epsffile[72 72 540 720]{"`zcat screendump.ps.Z"}

to include screendump.ps. Of course, the commands to be executed can be anything, including using a file conversion utility such as tek2ps or whatever is appropriate.

This extension is not portable to other dvi2ps programs. Yet.

5. Using PostScript Fonts

Thanks to Donald E. Knuth, the dvips driver now supports PostScript fonts through the virtual font capability. PostScript fonts are (or should be) accompanied by a font metric file such as Times-Roman.afm, which describes characteristics of a font called Times-Roman. To use such fonts with TEX, we need tfm files that contain similar information. These can be generated from afm files by running the program afm2tfm, supplied with dvips. This program also creates virtual fonts with which you can use normal plain TEX conventions.

Note that non-resident downloaded PostScript fonts tend to use a great deal of printer virtual memory. PostScript printers typically have between 130,000 and 400,000 bytes of available virtual memory; each downloaded font will require approximately 30,000 bytes of that. For many applications, bitmapped fonts work much better, even at typesetter resolutions (with the -Z option.)

Even resident PostScript fonts can take a fair amount of memory, but less with this release of dvips than previously. Also, bitmapped fonts tend to image faster than PostScript fonts.

A few Type 1 fonts (such as Utopia, Charter, and Courier) have been contributed by vendors to the X distribution, and are freely available. You can get TEX distributions for

them from ftp.cs.umb.edu in pub/tex, and from the CTAN hosts in tex-archive/fonts.

Your Unix system may have come with additional PostScript fonts. If so, you can make them available to Dvips by copying the files or making symbolic links with the appropriate filenames, and running afm2tfm to make TFM and VF files so the fonts will be available in the same encoding as the fonts distributed with dvips. Also check psfonts.map to be sure the fonts are listed there.

Here are the typical locations for vendor-supplied fonts:

           DEC Ultrix  /usr/lib/DPS/outline/decwin
           DEC OSF/1  /usr/lib/X11/fonts/Type1Adobe
           NeXT  /NextLibrary/Fonts/outline
           SGI IRIX  /usr/lib/DPS/outline/base
           Sun Solaris 2.3 /usr/openwin/lib/X11/fonts/Type1/outline

5.1 The afm2tfm Program

The afm2tfm program can convert an Adobe afm file into a ‘raw’ TEX tfm file with the command

    localhost> afm2tfm Times-Roman rptmr

(You should run this from in a directory where Times-Roman.afm resides.) The output file rptmr.tfm is ‘raw’ because it does no character remapping; it simply converts the character information on a one-to-one basis to TEX characters with the same code. The name rptmr stands for Resident PostScript Times Roman; section 6 below explains more about a proposed scheme for font names.

In the following examples, we will use the font Times-Roman to illustrate the conversion process. For the standard 35 LaserWriter fonts, however, it is highly recommended that you use the supplied tfm and vf files that come with dvips (usually in a file called dvipslib.tar.Z), as these files contain some additional changes that make them work better with TEX than they otherwise would.

PostScript fonts have a different encoding scheme from that of plain TEX. Although both schemes are based on ASCII, special characters such as ligatures and accents are handled quite differently. Therefore we obtain best results by using a ‘virtual font’ interface, which makes TEX act as if the PostScript font had a standard TEX encoding. Such a virtual font can be obtained, for example, by the command

    localhost> afm2tfm Times-Roman -v ptmr rptmr

This produces two outputs, namely the ‘virtual property list’ file ptmr.vpl, and the TEX font metric file rptmr.tfm. The latter file describes an ‘actual font’ on which the virtual font is based.

To use the font in TEX, you should first run

    localhost> vptovf ptmr.vpl ptmr.vf ptmr.tfm

and then install the virtual font file ptmr.vf in the virtual font directory (by default,

/usr/lib/tex/fonts/vf) and install ptmr.tfm and rptmr.tfm in the directory for TEX

font metrics (by default, /usr/lib/tex/fonts/tfm). (This probably has already been done for you for the most common PostScript fonts.) You can also make more complex virtual fonts by editing ptmr.vpl before running vptovf; such editing might add the uppercase Greek characters in the standard TEX positions, for instance. Once this has been done, you’re all set. You can use code like this in TEX henceforth:

    \font\myfont=ptmr at 10pt
    \myfont Hello, I am being typeset in Times-Roman.

Note that there are two fonts, one actual (‘rptmr’, which is analogous to a raw piece of hardware) and one virtual (‘ptmr’, which has typesetting know-how added). You could also say

    \font\TR=rptmr at 10pt

and typeset directly with that, but then you would have no ligatures or kerning, and you would have to use Adobe character positions for special letters like Æ. The virtual font called ptmr not only has ligatures and kerning, and most of the standard accent conventions of TEX, it also has a few additional features not present in the Computer Modern fonts. For example, it includes all the Adobe characters (such as the Polish ogonek and the French guillemots). The only things you lose from ordinary TEX text fonts are the dotless j (which can be hacked into the VPL file with literal PostScript specials if you have the patience) and uppercase Greek letters (which just don’t exist unless you buy them separately). Experts may refer to Donald E. Knuth article in TUGboat v. 11, no. 1, Apr. 1990, pp. 13–23. “Virtual Fonts: More Fun for Grand Wizards.”

When dvips goes to use a font, it first checks to see if it is one of the fonts listed in a file called psfonts.map. If it is mentioned in that file, then it is assumed that the font is a resident PostScript font and can be found with the PostScript findfont operator. This file resides by default in /usr/lib/tex/ps, and consists of a single font per line. Note that only the raw PostScript font names should be listed in this file; the vf fonts should not be, since they are automatically mapped to the raw PostScript fonts by the virtual font machinery. The supplied psfonts.map file defines the most common PostScript fonts.

As much as possible, the PostScript fonts follow plain TEX conventions for accents. The two exceptions to this are the Hungarian umlaut (which is at position 0x7D in cmr10, but position 0xCD in ptmr) and the dot accent (at positions 0x5F and 0xC7, respectively). In order to use these accents with PostScript fonts or in math mode when \textfont0 is a PostScript font, you will need to use the following definitions. Note that these definitions will not work with normal TEX fonts for the relevant accents; note also that these definitions are already part of the distributed psfonts.sty. In addition, the \AA that is used to typeset the Å character must be modified as shown.

    \def\H#1{{\accent"CD #1}}\def\.#1{{\accent"C7 #1}}
    \def\dot{\mathaccent"70C7 }
    \newdimen\aadimen
    \def\AA{\leavevmode\setbox0\hbox{h}\aadimen\ht0
      \advance\aadimen-1ex\setbox0\hbox{A}\rlap{\raise.67\aadimen
      \hbox to \wd0{\hss\char'27\hss}}A}

These PostScript fonts can be scaled to any size. Go wild! Note, however, that using PostScript fonts does use up a great deal of the printer’s virtual memory and it does take time. You may find downloading the Computer Modern bitmapped fonts to be faster than using the built-in PostScript fonts.

5.2 Changing a Font’s Encoding

The afm2tfm program also allows you to specify a different encoding for a PostScript font. This should only be done by wizards. This can be done at two levels.

You can specify a different output encoding with -o. This only applies when you are building a virtual font, and it tells afm2tfm to attempt to remap the font to match the output encoding as closely as possible. In such an output encoding, you can also specify ligature pairs and kerning information that will be used in addition to the information in the afm file. This will be the most common re-encoding required, since the only thing that changes is the vf file (and its associated tfm file) and since most everything you would want to do can be done with this method.

You can also specify a different PostScript encoding with -p. This option affects the generation of both the raw tfm file and the virtual vf and tfm files, and it also requires that the encoding file be available to be downloaded as part of every PostScript document produced that uses this font. But this is the only way to access characters in a PostScript font that are neither encoded in the afm file nor built from other characters (constructed characters.) For instance, Times-Roman contains the extra characters registered and thorn (among others) that can only be accessed through such a PostScript reencoding. Any ligature or kern information specified in the PostScript encoding is ignored by afm2tfm.

The format of the encoding files is simple—it is precisely the same format that is used to define an encoding vector in PostScript! For this reason, the same file can be used as a PostScript or TEX encoding file for afm2tfm as well as downloaded to the printer as part of a document that uses a reencoded font.

The specific format that afm2tfm accepts is one of the following form:

% comments are mostly ignored; more on this later
/MyEncoding [ /Alpha /Beta /Gamma /Delta ...
     /A /B ... /Z % exactly 256 entries, each with a / at the front
     /wfooaccent /xfooaccent /yfooaccent /zfooaccent ] def

Comments, which start with a percent sign and continue until the end of the current line, are mostly ignored. The first ‘word’ of the file must start with a forward slash (a PostScript literal name) and is used as the name of the encoding. The next word must be an open bracket. Following that must be precisely 256 character names; use /.notdef for any that you do not want to define. Finally, there must be a close bracket. The final token is usually def, but this is not enforced. Note that all names are case sensitive.

Any ligature or kern information is given in the comments. As each comment is encountered, it is examined. If the first word after the percent sign is LIGKERN, exactly, then the entire rest of the line is parsed for ligature and kern information. This ligature and kern information is given in groups of words, each group of which must be terminated by a semicolon (with a space before and after it, unless it occurs on the end of a line.)

In these LIGKERN statements, three types of information may be specified. These three types are ligature pairs, kerns to remove or ignore, and the character value of this font’s boundary character. Which of the types the particular set of words corresponds to is automatically determined by the allowable syntax.

Throughout the LIGKERN section, the boundary character is specified as ||. To set the boundary character value, a command such as || = 39 ; must be used.

To indicate a kern to remove, give the names of the two characters (without the leading slash) separated by {}, as in one {} one ;. This is similar to the way you might use {} in a TEX file to turn off ligatures or kerns at a particular location. Either or both of the character names can be given as *, which is a wild card matching any character; thus, all kerns can be removed with * {} * ;.

To specify a ligature, specify the names of the pair of characters, followed by the ligature ‘operation’ (as in ), followed by the replacing character name. Either (but not both) of the first two characters can be || to indicate a word boundary. Normally the ‘operation’ is =: meaning that both characters are removed and replaced by the third character, but by adding | characters on either side of the =:, you can specify which of the two leading characters to retain. In addition, by suffixing the ligature operation with one or two > signs, you can indicate that the ligature scanning operation should skip that many characters before proceeding. This works just like in . A typical ligature might be specified with ff i =: ffi ;. A more convoluted ligature is one one |=:|>> exclam ; which indicates that every pair of adjacent 1’s should be separated by an exclamation point, and then two of the resulting characters should be skipped over before continuing searching for ligatures and kerns. You cannot give more >’s in an ligature operation as you did |, so there are a total of eight possible ligature operations:

    =:  |=:  |=:> =:| =:|> |=:| |=:|> |=:|>>

The default set of ligatures and kerns built in to afm2tfm can be specified with:

% LIGKERN space l =: lslash ; space L =: Lslash ;
% LIGKERN question quoteleft =: questiondown ;
% LIGKERN exclam quoteleft =: exclamdown ;
% LIGKERN hyphen hyphen =: endash ; endash hyphen =: emdash ;
% LIGKERN quoteleft quoteleft =: quotedblleft ;
% LIGKERN quoteright quoteright =: quotedblright ;
% LIGKERN space {} * ; * {} space ; zero {} * ; * {} zero ;
% LIGKERN one {} * ; * {} one ; two {} * ; * {} two ;
% LIGKERN three {} * ; * {} three ; four {} * ; * {} four ;
% LIGKERN five {} * ; * {} five ; six {} * ; * {} six ;
% LIGKERN seven {} * ; * {} seven ; eight {} * ; * {} eight ;
% LIGKERN nine {} * ; * {} nine ;

5.3 Special Effects

Special effects are also obtainable, with commands such as

    localhost> afm2tfm Times-Roman -s .167 -v ptmro rptmro

which create ptmro.vpl and rptmro.tfm. To use this, proceed as above but put the line

    rptmro Times-Roman ".167 SlantFont"

into psfonts.map. Then rptmro (our name for an obliqued Times) will act as if it were a resident font, although it is actually constructed from Times-Roman by PostScript hackery. (This oblique version of Times-Roman is obtained by slanting everything 1/6 to the right.) Similarly, you can get an expanded font by

    localhost> afm2tfm Times-Roman -e 1.2 -v ptmrre rptmrre

and by recording the pseudo-resident font

    rptmrre Times-Roman "1.2 ExtendFont"
in psfonts.map.

You can also create a small caps font with a command such as

    localhost> afm2tfm Times-Roman -V ptmrc rptmr

This is done strictly with a virtual font, however. In addition, the font on which the small caps font is based (in this case rptmr may already be created and installed, in which case no additional psfonts.map entry is needed. In any case, you must give the appropriate name of the font that is not small caps as the base name (last parameter) to afm2tfm. For instance, if you create a slanted small caps font, you must give the base name of the raw slanted font as that last parameter, not the base name of the unslanted font.

By default, the -V option uses a font scaled to 80% for lower case. If you specify the -c option, you can change this scaling.

If you change the PostScript encoding of a font, you must specify the input file as a header file, as well as give a reencoding command. For instance, let us say we are using Times-Roman reencoded according to the encoding MyEncoding (stored in the file myenc.enc) as rptmrx. In this case, our psfonts.map entry would look like

    rptmrx Times-Roman "MyEncoding ReEncodeFont" <myenc.enc

The afm2tfm program prints out the precise line you need to add to psfonts.map to use that font, assuming the font is resident in the printer; if the font is not resident, you must add the header command to download the font yourself. Note that each identical line only needs to be specified once in the psfonts.map file, even though many different fonts (small caps variants, or ones with different output encodings) may be based on it.

The command line switches to afm2tfm are:

-e ratio All characters are stretched horizontally by the stated ratio; if it is less than 1.0, you get a condensed font.

-c scale If this option is given when creating a small caps font (with -V), then the scaling for the ‘lower’ case will be changed from the default 0.8 to the fraction given here.

-O This option forces all character designations in the resultant vpl file be given as octal values; this is useful for symbol or other special-purpose fonts where character names such as ‘A’ have no meaning.

-p file This specifies a file to use for the PostScript encoding of the font. Note that this file must also be mentioned as a header file for the font in psfonts.map, and that ligature and kern information in this file is ignored.

-s slant All characters are slanted to the right by the stated slant; if it is negative, the letters slope to the left (or they might be upright if you start with an italic font).

-t file This specifies a file to use for the target TEX encoding of the font. Ligature and kern information may also be specified in this file; the file is not needed once the vf file has been created.

-T file This option specifies that file is to be used for both the PostScript and target TEX encodings of the font.

-u This option indicates that afm2tfm should use only those characters that are required by the output encoding, and no others. Normally, afm2tfm tries to include both characters that fit the output encoding and any additional characters that might exist in the font. This option forbids those additional characters from being added.

-v file Generate a virtual property list vpl file as well as a tfm file.

-V file Same as -v, but the virtual font generated is a small caps font obtained by scaling uppercase letters by 0.8 to typeset lowercase. This font handles accented letters and retains proper kerning.

5.4 Non-Resident PostScript Fonts

If you want to use a non-printer-resident PostScript font for which you have a pfb or pfa file (an Adobe Type 1 font program), you can make it act like a resident font by putting a ‘<’ sign and the name of the pfb or pfa file just after the font name in the psfonts.map file entry. For example,

    rpstrn StoneInformal <StoneInformal.pfb

will cause dvips to include StoneInformal.pfb in your document as if it were a header file, whenever the pseudo-resident font StoneInformal is used in a document. Similarly, you can generate transformed fonts and include lines like

    rpstrc StoneInformal <StoneInformal.pfb ".8 ExtendFont"

in psfonts.map, in which case StoneInformal.pfb will be loaded whenever StoneInformal-Condensed is used. (Each header file is loaded at most once per PostScript file. The pfb files should be installed in the dvips header directory [usually /usr/lib/tex/ps] with the other header files.)

If you are using a pfb file that has different PostScript encodings, you would need to multiple header files for that font in psfonts.map. If, for instance, StoneInformal was both non-resident and you wanted to reencode it in PostScript with MyEncoding stored in myenc.enc, a line such as

rpstrnx StoneInformal "MyEncoding ReEncodeFont" <myenc.enc <StoneInformal.pfb

When using such files, dvips is smart enough to unpack the standard binary pfb format into ASCII so there is no need to perform this conversion yourself. In addition, it will scan the font to determine its memory usage, as it would for any header file.

5.5 Font Aliases

Some systems don’t handle files with long names well—MS-DOS is a notable example. For this reason, dvips will accept an alias for such fonts. Such an alias should be the first word in the psfonts.map line. For instance, if we wanted the name rptmr to be used for the raw Times-Roman since our computer can’t handle long names or, alternatively, we want to follow the standard naming conventions, we would use the following line in our

psfonts.map file:
    rptmr Times-Roman

The tfm file must have the name rptmr.tfm.

The distribution file adobe contains a list of the short names that should be used for most Adobe fonts currently available. Please reference this file when installing a new font and use the standard name.

The parsing of the psfonts.map file should be explained to eliminate all confusion. If a line is empty or begins with a space, asterisk, semicolon, or hash mark, it is ignored. Each remaining line is separated into words, where words are separated by spaces or tabs. If a word begins with a double quote, it extends until the next double quote or the end of the line. If a word starts with a less than character, it is treated as a font header file (or a downloaded PostScript font). There can be more than one such header for a given font. If a word starts with a double quote, it is a special instruction for generating that font. Otherwise it is a name. The first such name is always the name TEX uses for the font and is also the name of the raw tfm file. If there is another name word, that name is used as the PostScript name; if there is only one name word, it is used for both the TEX name and the PostScript name.

Note that dvips no longer registers the full PostScript name if an alias is given, so the single line

rptmr Times-Roman

would only allow dvips to find the rptmr font and not the Times-Roman font.

6. Font Naming Conventions

This section of the manual has been written by Karl Berry and specifies a standard for naming fonts for TEX. This standard has been adopted in dvips, and it is recommended that it be followed where possible.

As more typeface families become available for use with TEX, the need for a consistent, rational naming scheme for the font filenames concomitantly grows. Some (electronic) discussion has gone into the following proposal; I felt it was appropriate now to bring it before a wider community. In some respects, it follows and simplifies Mittelbach and Schöpf’s article in TUGboat, volume 11, number 2 (June 1990).

Here are some facts about fonts that went into the hopper when creating this proposal:

• TEX runs on virtually all computers, under almost as many operating systems, all with their own idea of how files should be named. Any proposal regarding filenames, therefore, must cater to the lowest common denominator. That seems to be eight characters in length, not counting any extension, and with case being insignificant. Characters other than letters and numerals are probably unusable.

• Most typefaces are offered by several vendors. The version offered by vendor A is not compatible with that of vendor B.

• Typefaces typically come in different weights (hairline to extra heavy), different expansions (ultra condensed to wide), and an open-ended range of variants (italic, sans serif, typewriter, shadow, …). No accepted standards exist for any of these qualities, nor are any standards ever likely to gain acceptance.

• The Computer Modern typeface family preserves traditional typesetting practice in at least one important respect: different sizes of the same font are not scaled linearly. This is in contrast to most commercial fonts available.

Here is how I propose to divide up the eight characters:

FTTWVEDD

where • F represents the foundry that produced the font, and is omitted if there isn’t one.

• TT represents the typeface name.

• W represents the weight.

• V represents the variant, and is omitted if both it and the expansion are “normal”.

• E represents the expansion, and is omitted if it is “normal”.

• DD represents the design size, and is omitted if the font is linearly scaled from a single tfm file.

See the section on virtual fonts (towards the end) for an exception to the above.

The weight, variant, and expansion are probably all best taken from the original source of the typeface, instead of trying to relate them to some external standard.

Before giving the lists of abbreviations, let me point out two problems, to neither of which I have a good solution. 1) Assuming that only the English letters are used, two letters is enough for only 676 typeface families (even assuming we want to use all possible combinations, which is doubtful). There are many more than 676 typeface families in the world. 2) Fonts with design sizes over 100 pt are not common, but neither are they unheard of.

On to the specifics of the lists. If you adopt this proposal at your own installation, and find that you have fonts with some property I missed, please write to me (see the end of the article for various addresses), so I can update the lists. You can get the most up-to-date version of these lists electronically, by anonymous ftp from the host ftp.cs.umb.edu. I will also send them to you by electronic mail, if necessary.

I give the letters in lowercase, which is preferred on systems where case is significant. Most lists are in alphabetical order by the abbreviations.

6.1 Foundry

This is the current list of foundries.

DVIPS: A TEX Driver 22

a Autologic
b Bitstream
c Agfa-Compugraphic
g Free Software Foundation (g for GNU)
h Bigelow & Holmes (with apologies to Chuck)
i International Typeface Corporation
p Adobe (p for PostScript)
r reserved for use with virtual fonts; see below
s Sun

6.2 Typeface Families

The list of typefaces is:

ad Adobe Garamond  go Goudy Oldstyle
ag Avant Garde  gs Gill Sans
ao Antique Olive  jo Joanna
at American Typewriter  lc Lucida
bb Bembo  lt Lutetia
bd Bodoni  nc New Century Schoolbook
bg Benguiat  op Optima
bk Bookman  pl Palatino
bl Balloon  pp Perpetua
bv Baskerville  rw Rockwell
bw Broadway  st Stone
cb Cooper Black  sy Symbol
cl Cloister  tm Times
cr Courier  un Univers
cn Century  uy University
cs Century Schoolbook  zc Zapf Chancery
hv Helvetica  zd Zapf Dingbats
gm Garamond

6.3 Weight

This is a list of the possible weights, roughly in order from lightest to heaviest.

a hairline  d demi
t thin  s semi
i extra light  b bold
l light  x extra bold
k book  h heavy
r regular  c black
m medium  u ultra

6.4 Variants

The variants are:

a alternate  n informal
b bright  o oblique (i.e., slanted)
c small caps  r normal (roman or sans)
e engraved  s sans serif
g grooved (as in the IBM logo)  t typewriter
h shadow  u unslanted italic
i (text) italic  x expert
l outline

If the variant is r, and the expansion is also normal, both the variant and the expansion are omitted. When the normal version of the typeface is sans serif (e.g., Helvetica), r should be used, not s. Use s only when the typeface family has both serif and sans serif variants. The “alternate” variant (a) is used by some Adobe fonts that have spiffy swashes and additional ligatures. The “expert” variant (x) is also used by some Adobe fonts with oldstyle figures and small caps.

Some fonts have multiple variants; Stone Informal Italic, for example. The only reasonable approach to these is to list all the letters for all the variants, choosing one to end with that is not also an expansion letter. Of course, it is possible that the name will become too long if you do this, but … well, I’m open to suggestions. It’s also possible the name will be ambiguous, if some new letter is used for expansions in the future. You can avoid this problem by adding the expansion r (if it doesn’t make the name too long), and never using r for the last variant.

6.5 Expansion

This is a list of the possible expansions, in order from narrowest to widest.

o extra condensed  x extended (by hand)
c condensed (by hand)  e expanded (automatic)
n narrow (automatic)  w wide
r regular, normal, medium (usually omitted)

Expansion of fonts is sometimes done automatically (as in PostScript scale), and sometimes done by humans. I chose ‘narrow’ and ‘expanded’ to imply the former, and ‘condensed’ and ‘extended’ to imply the latter, as I believe this reflects common usage.

6.6 Naming Virtual Fonts

In concert with releasing TEX 3.0 and 2.0, Don Knuth wrote two new utility programs: VFtoVP and VPtoVF, which convert to and from “virtual” fonts. Virtual fonts provide a general interface between the writers of TEX macros and font suppliers. In general, therefore, it is impossible to come up with a general scheme for naming virtual fonts, since each virtual font is an individual creation, possibly bringing together many unrelated fonts.

Nevertheless, one common case is to use virtual fonts to map TEX’s default accent and other character code conventions onto a vendor-supplied font. For example, dvips does this for fonts given in the PostScript “standard encoding”. In this case, each font consists of a “virtual” tfm file, which is what TEX uses, a “raw” tfm file, which corresponds to the actual device font, and a vf file, which describes the relationship between the two.

This adds another dimension to the space of font names, namely, “virtualness” (or rather, “rawness”, since it is the virtual tfm files that the users want to see). But we have already used up all eight characters in the font names.

The best solution I have been able to think of is this: prepend r to the raw tfm files; the virtual tfm files should be named with the usual foundry prefix. For example, the virtual Times-Roman tfm file is named ptmr, as usual; the raw Times-Roman tfm file is named rptmr. To prevent intolerable confusion, I promise never to give a foundry the letter r.

This scheme will work only as long as the virtualized fonts do not have design sizes; if they do, another foundry letter will have to be allocated, it seems to me.

A pox upon the houses of those who decided on fixed-length filenames!

6.7 Examples

In closing, I will give two examples. First, the fonts in the Univers typeface family were assigned numbers by its designer, Adrien Frutiger. (You can see the scheme on, for example, page 29 of The Art of Typo.icon.ography, by Martin Solomon.) Naturally, we want to give them names.

unl 45 (light)  unmro 59 (medium extra condensed)
unli 46 (light italic)  undrx 63 (demibold extended)
unlrc 47 (light condensed)  und 65 (demibold)
unlic 48 (light condensed italic)  undi 66 (demibold italic)
unlro 49 (light extra condensed)  undrc 67 (demibold condensed)
unmrx 53 (medium extended)  undic 68 (demibold condensed italic)
unm 55 (medium)  unbrx 73 (bold extended)
unmi 56 (medium italic)  unb 75 (bold)
unmrc 57 (medium condensed)  unbi 76 (bold italic)
unmic 58 (medium condensed italic)  unxrx 83 (extra bold extended)

Second, here are names for the standard PostScript fonts and their variants: Fonts marked by an asterisk do not require using virtual fonts; the raw fonts can be used directly because no remapping is necessary; every character is encoded.

pagk AvantGarde-Book  pncri NewCenturySchlbk-Italic
pagkc AvantGarde-Book (Small Caps)  pncr NewCenturySchlbk
pagko AvantGarde-BookOblique  pncrc NewCenturySchlbk (Small Caps)
pagd AvantGarde-Demi  pplb Palatino-Bold
pagdo AvantGarde-DemiOblique  pplbi Palatino-BoldItalic
pbkd Bookman-Demi  pplbu Palatino-BoldUnslanted
pbkdi Bookman-DemiItalic  pplrrn Palatino-Narrow
pbkl Bookman-Light  pplrre Palatino-Expanded
pbkli Bookman-LightItalic  pplri Palatino-Italic
pbklc Bookman-Light (Small Caps)  pplr Palatino
pcrb Courier-Bold  pplro Palatino-Oblique
pcrbo Courier-BoldOblique  pplru Palatino-Unslanted
pcrro Courier-Oblique  pplrc Palatino (Small Caps)
pcrr Courier  psyr Symbol^{∗}
phvb Helvetica-Bold  psyro Symbol-Oblique^{∗}
phvbo Helvetica-BoldOblique  ptmb Times-Bold
phvro Helvetica-Oblique  ptmbi Times-BoldItalic
phvr Helvetica  ptmrrn Times-Narrow
phvrc Helvetica (Small Caps)  ptmrre Times-Expanded
phvbrn Helvetica-Narrow-Bold  ptmri Times-Italic
phvbon Helvetica-Narrow-BoldOblique  ptmro Times-Oblique
phvron Helvetica-Narrow-Oblique  ptmr Times-Roman
phvrrn Helvetica-Narrow  ptmrc Times-Roman (Small Caps)
pncb NewCenturySchlbk-Bold  pzcmi ZapfChancery-MediumItalic
pncbi NewCenturySchlbk-BoldItalic  pzdr ZapfDingbats^{∗}

Please contact Karl Berry if you have any comments or additions. Karl can be reached at karl@cs.umb.edu, or at 135 Center Hill Road, Plymouth, MA 02360.

7. Command Line Options

The dvips driver has a plethora of command line options. Reading through this section will give a good idea of the capabilities of the driver.

Many of the parameterless options listed here can be turned off by immediately suffixing the option with a zero (0); for instance, to turn off page reversal if it is turned on by default, use -r0. The options that can be turned off in this way are a, f, k, i, m, q, r, s, E, F, K, M, N, U, and Z.

This is a handy summary of the options; it is printed out when you run dvips with no arguments.

    Usage: dvips [options] filename[.dvi]
a*  Conserve memory, not time      y # Multiply by dvi magnification
b # Page copies, for posters e.g.  A   Print only odd (TeX) pages
c # Uncollated copies              B   Print only even (TeX) pages
d # Debugging                      C # Collated copies
e # Maxdrift value                 D # Resolution
f*  Run as filter                  E*  Try to create EPSF
h f Add header file                F*  Send control-D at end
i*  Separate file per section      K*  Pull comments from inclusions
k*  Print crop marks               M*  Don't make fonts
l # Last page                      N*  No structured comments
m*  Manual feed                    O c Set/change paper offset
n # Maximum number of pages        P s Load config.$s
o f Output file                    R   Run securely
p # First page                     S # Max section size in pages
q*  Run quietly                    T c Specify desired page size
r*  Reverse order of pages         U*  Disable string param trick
s*  Enclose output in save/restore V*  Send downloadable PS fonts as PK
t s Paper format                   X # Horizontal resolution
x # Override dvi magnification     Y # Vertical resolution
                                   Z*  Compress bitmap fonts
    # = number   f = file   s = string  * = suffix, `0' to turn off
    c = comma-separated dimension pair (e.g., 3.2in,-32.1cm)

-a: Conserve memory by making three passes over the dvi file instead of two and only loading those characters actually used. Generally only useful on machines with a very limited amount of memory, like some PCs.

-b num: Generate num copies of each page, but duplicating the page body rather than using the #numcopies option. This can be useful in conjunction with a header file setting \bop-hook to do color separations or other neat tricks.

-c num: Generate num copies of every page, by using PostScript’s #copies feature. Default is 1. (For collated copies, see the -C option below.)

-d num: Set the debug flags. This is intended only for emergencies or for unusual factfinding expeditions; it will work only if dvips has been compiled with the DEBUG option. The source file debug.h indicates what the values of num can be, or see section 15 of this manual. Use a value of −1 for maximum output.

-e num: Make sure that each character is placed at most this many pixels from its ‘true’ resolution-independent position on the page. The default value of this parameter is resolution dependent (it is the number of entries in the list [100, 200, 300, 400, 500, 600, 800, 1000, 1200, 1600, 2000, 2400, 2800, 3200, …] that are less than or equal to the resolution in dots per inch). Allowing individual characters to ‘drift’ from their correctly rounded positions by a few pixels, while regaining the true position at the beginning of each new word, improves the spacing of letters in words.

-f: Run as a filter. Read the dvi file from standard input and write the PostScript to standard output. The standard input must be seekable, so it cannot be a pipe. If you must use a pipe, write a shell script that copies the pipe output to a temporary file and then points dvips at this file. This option also disables the automatic reading of the PRINTER environment variable, and turns off the automatic sending of control D if it was turned on with the -F option or in the configuration file; use -F after this option if you want both.

-h name: Prepend file name as an additional header file. (However, if the name is simply ‘-’, suppress all header files from the output.) This header file gets added to the PostScript userdict.

-i: Make each section be a separate file. Under certain circumstances, dvips will split the document up into ‘sections’ to be processed independently; this is most often done for memory reasons. Using this option tells dvips to place each section into a separate file; the new file names are created replacing the suffix of the supplied output file name by a three-digit sequence number. This option is most often used in conjunction with the -S option which sets the maximum section length in pages. For instance, some phototypesetters cannot print more than ten or so consecutive pages before running out of steam; these options can be used to automatically split a book into ten-page sections, each to its own file.

-k: Print crop marks. This option increases the paper size (which should be specified, either with a paper size special or with the -T option) by a half inch in each dimension. It translates each page by a quarter inch and draws cross-style crop marks. It is mostly useful with typesetters that can set the page size automatically.

-l num: The last page printed will be the first one numbered num. Default is the last page in the document. If the num is prefixed by an equals sign, then it (and any argument to the -p option) is treated as a sequence number, rather than a value to compare with \count0 values. Thus, using -l =9 will end with the ninth page of the document, no matter what the pages are actually numbered.

-m: Specify manual feed for printer.

-n num: At most num pages will be printed. Default is 100000.

-o name: The output will be sent to file name. If no file name is given, the default name is file.ps where the dvi file was called file.dvi; if this option isn’t given, any default in the configuration file is used. If the first character of the supplied output file name is an exclamation mark, then the remainder will be used as an argument to popen; thus, specifying !lpr as the output file will automatically queue the file for printing. This option also disables the automatic reading of the PRINTER environment variable, and turns off the automatic sending of control D if it was turned on with the -F option or in the configuration file; use -F after this option if you want both.

-p num: The first page printed will be the first one numbered num. Default is the first page in the document. If the num is prefixed by an equals sign, then it (and any argument to the -l option) is treated as a sequence number, rather than a value to compare with \count0 values. Thus, using -p =3 will start with the third page of the document, no matter what the pages are actually numbered. Another form of page selection is available by using -pp followed by a comma-separated list of pages or page-ranges, where the page ranges are colon-separated pairs of numbers. Thus, you can print pages 3–10, 21, and 73–92 with the option -pp 3:10,21,73:92.

-q: Run in quiet mode. Don’t chatter about pages converted, etc.; report nothing but errors to standard error.

-r: Stack pages in reverse order. Normally, page 1 will be printed first.

-s: Causes the entire global output to be enclosed in a save/restore pair. This causes the file to not be truly conformant, and is thus not recommended, but is useful if you are driving the printer directly and don’t care too much about the portability of the output.

-t papertype: This sets the paper type to papertype. The papertype should be defined in one of the configuration files, along with the appropriate code to select it. See the documentation for @ in the configuration file option descriptions. You can also specify -t landscape, which rotates a document by 90 degrees. To rotate a document whose size is not letter, you can use the -t option twice, once for the page size, and once for landscape. The upper left corner of each page in the dvi file is placed one inch from the left and one inch from the top. Use of this option is highly dependent on the configuration file. Note that executing the letter or a4 or other PostScript operators cause the document to be nonconforming and can cause it not to print on certain printers, so the default paper size should not execute such an operator if at all possible.

-x num: Set the magnification ratio to num/1000. Overrides the magnification specified in the dvi file. Must be between 10 and 100000. It is recommended that you use standard magstep values (1095, 1200, 1440, 1728, 2074, 2488, 2986, and so on) to help reduce the total number of PK files generated.

-A: This option prints only the odd pages. This option uses the TEX page numbering rather than the sequence page numbers.

-B: This option prints only the even pages. This option uses the TEX page numbering rather than the sequence page numbers.

-C num: Create num copies, but collated (by replicating the data in the PostScript file). Slower than the -c option, but easier on the hands, and faster than resubmitting the same PostScript file multiple times.

-D num: Set the resolution in dpi (dots per inch) to num. This affects the choice of bitmap fonts that are loaded and also the positioning of letters in resident PostScript fonts. Must be between 10 and 10000. This affects both the horizontal and vertical resolution. If a high resolution (something greater than 400 dpi, say) is selected, the -Z flag should probably also be used.

-E: Makes dvips attempt to generate an EPSF file with a tight bounding box. This only works on one-page files, and it only looks at marks made by characters and rules, not by any included graphics. In addition, it gets the glyph metrics from the tfm file, so characters that lie outside their enclosing tfm box may confuse it. In addition, the bounding box might be a bit too loose if the character glyph has significant left or right side bearings. Nonetheless, this option works well for creating small EPSF files for equations or tables or the like. (Note, of course, that dvips output is resolution dependent and thus does not make very good EPSF files, especially if the images are to be scaled; use these EPSF files with a great deal of care.)

-F: Causes Control-D (ASCII code 4) to be appended as the very last character of the PostScript file. This is useful when dvips is driving the printer directly instead of working through a spooler, as is common on extremely small systems. Otherwise, it is not recommended.

-K: This option causes comments in included PostScript graphics, font files, and headers to be removed. This is sometimes necessary to get around bugs in spoolers or PostScript post-processing programs. Specifically, the %%Page comments, when left in, often cause difficulties. Use of this flag can cause some included graphics to fail, since the PostScript header macros from some software packages read portions of the input stream line by line, searching for a particular comment. This option has been turned on by default because PostScript previewers and spoolers still have problems with the structuring conventions.

-M: Turns off the automatic font generation facility. If any fonts are missing, commands to generate the fonts are appended to the file missfont.log in the current directory; this file can then be executed and deleted to create the missing fonts.

-N: Turns off structured comments; this might be necessary on some systems that try to interpret PostScript comments in weird ways, or on some PostScript printers. Old versions of TranScript in particular cannot handle modern Encapsulated PostScript.

-O offset: Move the origin by a certain amount. The offset is a comma-separated pair of dimensions, such as .1in,-.3cm (in the same syntax used in the papersize special). The origin of the page is shifted from the default position (of one inch down, one inch to the right from the upper left corner of the paper) by this amount.

-P printername: Sets up the output for the appropriate printer. This is implemented by reading in config.printername, which can then set the output pipe (as in, o !lpr -Pprintername) as well as the font paths and any other defaults for that printer only. It is recommended that all standard defaults go in the one master config.ps file and only things that vary printer to printer go in the config.printername files. Note that config.ps is read before config.printername. In addition, another file called ~/.dvipsrc is searched for immediately after config.ps; this file is intended for user defaults. If no -P command is given, the environment variable PRINTER is checked. If that variable exists, and a corresponding configuration file exists, that configuration file is read in.

-S num: Set the maximum number of pages in each ‘section’. This option is most commonly used with the -i option; see that documentation above for more information.

-T offset: Set the paper size to the given pair of dimensions. This option takes its arguments in the same style as -O. It overrides any paper size special in the dvi file.

-U: Disable a PostScript virtual memory saving optimization that stores the character metric information in the same string that is used to store the bitmap information. This is only necessary when driving the Xerox 4045 PostScript interpreter. It is caused by a bug in that interpreter that results in ‘garbage’ on the bottom of each character. Not recommended unless you must drive this printer.

-V: Download non-resident PostScript fonts as bitmaps. This requires use of mtpk or pstopk or some combination of the two in order to generate the required bitmap fonts; neither of these programs are supplied with dvips.

-X num: Set the horizontal resolution in dots per inch to num.

-Y num: Set the vertical resolution in dots per inch to num.

-Z: Causes bitmapped fonts to be compressed before they are downloaded, thereby reducing the size of the PostScript font-downloading information. Especially useful at high resolutions or when very large fonts are used. Will slow down printing somewhat, especially on early 68000-based PostScript printers.

8. Configuration File Searching

The dvips program has a system of loading configuration files such that certain parameters can be set globally across the system, others can be set on a per-printer basis, and yet others can be set by the user. When dvips starts up, first the global config.ps file is searched for and loaded. This file is looked for along the path for configuration files, which is by default .:/usr/lib/tex/ps. After this master configuration file is loaded, a file by the name of .dvipsrc is loaded from the current user’s home directory, if such a file exists. This file is loaded in exactly the same way as the global configuration file, and it can override any options set in the global file.

Then the command line is read and parsed. If the -P option is encountered, at that point in the command line a configuration file for that printer is read in. Thus, the printer configuration file can override anything in the global or user configuration file, and it can also override anything seen in the command line up to the point that the -P option was encountered.

After the command line has been completely scanned, if there was no -P option selected, and also the -o and -f command line options were not used, a PRINTER environment variable is searched for. If this variable exists, and a configuration file for the printer mentioned in it exists, this configuration file is loaded last of all.

Note that because the printer-specific configuration files are read after the user’s configuration file, the user’s .dvipsrc cannot override things in the printer configuration files. On the other hand, the configuration path usually includes the current directory, and can be set to include the user’s home directory (or any other directory of the user), so the user can always provide personalized printer-specific configuration files that will be found before the system global ones.

If your printer uses a different resolution than 300 dpi, make sure that you have given a mode as well as a resolution in the printer configuration file. Also make sure that knows about the mode, by entering it into your local mode def file (typically waits.mf; amiga.mf on the Amiga, next.mf on the NeXT) and recreating the plain.base file for , including the mode def file. (Another good mode definition file is

modes.mf by Karl Berry, which is available from ftp.cs.umb.edu in pub/tex/modes.mf.)

The most common problem in generating fonts with is that this file with the mode definitions is not included when creating the plain.base file.

9. Configuration File Options

Most of the configuration file options are similar to the command line options, but there are a few new ones.

Again, many may be turned off by suffixing the letter with a zero (0). These options are a, f, q, r, I, K, N, U, and Z.

Within a configuration file, any empty line or line starting with a space, asterisk, equal sign, or a pound sign is ignored.

@ name hsize vsize: This option is used to set the paper size defaults and options for the particular printer this configuration file describes. There are three formats for this option. If the option is specified on a line by itself, with no parameters, it instructs dvips to discard all other paper size information (possibly from another configuration file) and start fresh. If three parameters are given, as above, with the first parameter being a name and the second and third being a dimension (as in 8.5in or 3.2cc, just like in the papersize special), then the option is interpreted as starting a new paper size description, where name is the name and hsize and vsize describe the horizontal and vertical size of the sheet of paper, respectively. If both hsize and vsize are equal to zero (although you must still specify units!) then any page size will match it. If the @ character is immediately followed by a + sign, then the remainder of the line (after skipping any leading blanks) is treated as PostScript code to send to the printer to select that particular paper size. After all that, if the first character of the line is an exclamation point, then the line is put in the initial comments section of the final output file; else, it is put in the setup section of the output file. For instance, a subset of the paper size information supplied in the default config.ps looks like

  @ letterSize 8.5in 11in
  @ letter 8.5in 11in
  @+ %%BeginPaperSize:  Letter
  @+ letter
  @+ %%EndPaperSize
  @ legal 8.5in 14in
  @+ !  %%DocumentPaperSizes:  Legal
  @+ %%BeginPaperSize:  Legal
  @+ legal
  @+ %%EndPaperSize

Note that you can even include structured comments in the configuration file! When dvips has a paper format name given on the command line, it looks for a match by the name; when it has a papersize special, it looks for a match by dimensions. The first match found (in the order the paper size information is found in the configuration file) is used. If nothing matches, a warning is printed and the first paper size given is used, so the first paper size should always be the default. The dimensions must match within a quarter of an inch. Landscape mode for all of the paper sizes are automatically supported. If your printer has a command to set a special paper size, then give dimensions of 0in 0in; the PostScript code that sets the paper size can refer to the dimensions the user requested as \hsize and \vsize; these will be macros defined in the PostScript that return the requested size in default PostScript units. Note that virtually all of the PostScript commands you use here are device dependent and degrade the portability of the file; that is why the default first paper size entry should not send any PostScript commands down (although a structured comment or two would be okay). Also, some printers want BeginPaperSize comments and paper size setting commands; others (such as the NeXT) want PaperSize comments and they will handle setting the paper size. There is no solution I could find that works for both (except maybe specifying both). See the supplied config.ps file for a more realistic example.

a: Conserve memory by making three passes over the dvi file instead of two and only loading those characters actually used. Generally only useful on machines with a very limited amount of memory, like some PCs. b num: Generate num copies of each page, but duplicating the page body rather than using

PostScript’s #copies. This can be useful in conjunction with a header file setting \bop-hook to do color separations or other neat tricks.

e num: Set the maximum drift parameter to num dots (pixels) as explained above.

f: Run as a filter by default.

h name: Add name as a PostScript header file to be downloaded at the beginning.

i num: Make each section be a separate file, and set the maximum number of pages in a given file to num. Under certain circumstances, dvips will split the document into ‘sections’ to be processed independently; this is most often done for memory reasons. Using this option tells dvips to place each section into a separate file; the new file names are created by replacing the suffix of the supplied output file name with a three-digit sequence number. This is essentially a combination of the command line options -i and -S; see the documentation for these options for more information.

m num: The value num is the virtual memory available for fonts and strings in the printer. Default is 180000. This value must be accurate if memory conservation and document splitting is to work correctly. To determine this value, send the following file to the printer:

  %!  Hey, we're PostScript
  /Times-Roman findfont 30 scalefont setfont 144 432 moveto
  vmstatus exch sub 40 string cvs show pop showpage

Note that the number returned by this file is the total memory free; it is often a good idea to tell dvips that the printer has somewhat less memory. This is because many programs download permanent macros that can reduce the memory in the printer. In general, a memory size of about 300000 is plenty, since dvips can automatically split a document if required. It is unfortunate that PostScript printers with much less virtual memory still exist. Some systems or printers can dynamically increase the memory available to a PostScript interpreter, in which case this file might return a ridiculously low number; the NeXT computer is such a machine. For these systems, a value of one million works well.

o name: The default output file is set to name. As above, it can be a pipe. Useful in printerspecific configuration files to redirect the output to a particular printer queue.

p name: The file to examine for PostScript font aliases is name. It defaults to psfonts.map. This option allows different printers to use different resident fonts. If the name starts with a ‘+’ character, then the rest of the name (after any leading spaces) is used as an additional map file; thus, it is possible to have local map files pointed to by local configuration files that append to the global map file. q: Run in quiet mode by default.

r: Reverse the order of pages by default.

s: Enclose the entire document in a global save/restore pair by default. Not recommended, but useful in some environments; this breaks the conformance of the document to the Adobe PostScript structuring conventions.

D num: Set the vertical and horizontal resolution to num dots per inch. Useful in printerspecific configuration files.

E command: Execute the system command listed, for example as a UNIX shell command. Execution takes place immediately, while the configuration file is being read. This option can be used to insert the current date into a header file, for instance, as explained at the end of section 13. Possibly dangerous; in many installations it may be disabled, in which case a warning message will be printed if the option is used.

H path: The (colon-separated) path to search for PostScript header files is path.

I: Ignore the PRINTER environment variable.

K: Filter comments out of included PostScript files; see the description above for more information.

M mode: Set mode as the mode to be used when generating fonts. This is passed along to MakeTeXPK and overrides mode derivation from the base resolution.

N: Disable PostScript comments by default.

O offset: Move the origin by a certain amount. The offset is a comma-separated pair of dimensions, such as .1in,-.3cm (in the same syntax as used in the papersize special). The origin of the page is shifted from the default position (of one inch down, one inch to the right from the upper left corner of the paper) by this amount. This is useful for a printer that consistently offsets output pages by a certain amount. You can use the file testpage.tex to determine the correct value for your printer. Be sure to do several runs with the same O value—some printers vary widely from run to run.

P path: The (colon-separated) path to search for bitmap pk font files is path. The TEXPKS environment variable will override this. If a % character is found in path, the following substitutions will be made, and then a search will be made for the resulting filenames. A %f is replaced by the font name. A %b is replaced by the output device horizontal resolution dots per inch. A %d is replaced by the font size in dots per inch. A %p is replaced by the font family; this is always pk. A %m is replaced by the font mode; this is the mode given in the M option. Note that, for each path element, if it contains a percent sign, you must give the full file name, including path, rather than just the directory name; a path element such as /fonts/%b will try to open /fonts/300 when looking for cmr10.329pk, for instance, and this may not be what is intended; /fonts/%b/%f.%dpk is needed. If a path element does not contain a percent sign, there is no need to specify the entire file name (because you can’t, unless you list all possible specific font names!).

R num num …: Sets up a list of default resolutions to search for pk fonts, if the requested size is not available. The output will then scale the font found using PostScript scaling to the requested size. Note that the resultant output will be ugly, and thus a warning is issued. To turn this off, use a line with just the R and no numbers.

S path: The path to search for special illustrations (Encapsulated PostScript files or psfiles) is path. The TEXINPUTS environment variable will override this.

T path: The path to search for the tfm files is path. The TEXFONTS environment variable will override this. This path is used for resident fonts and fonts that can’t otherwise be found. It’s usually best to make it identical to the path used by TEX.

U: Turns off a memory-saving optimization; this is necessary for the Xerox 4045 printer, but not recommended otherwise. See the description above for more information.

V path: The path to search for virtual font vf files is path. This may be device-dependent if you use virtual fonts to simulate actual fonts on different devices.

W string: Sends string to stderr, if a parameter is given; otherwise it cancels another previous message. This is useful in the default configuration file if you want to require the user to specify a printer, for instance, or if you want to notify the user that the resultant output has special characteristics.

X num: Set the horizontal resolution to num dots per inch.

Y num: Set the vertical resolution to num dots per inch.

Z: Compress all downloaded fonts by default, as above.

10. Automatic Font Generation

One major problem with TEX and the Computer Modern fonts is the huge amount of disk space a full set of high resolution fonts can take. The dvips program solves this problem by creating fonts on demand, so only those fonts that are actually used are stored on disk. At a typical site, less than one-fifth of the full set of Computer Modern fonts are used over a long period, so this saves a great deal of disk space.

Furthermore, the addition of dynamic font generation allows fonts to be used at any size, including typesetter resolutions and extremely huge banner sizes. Nothing special needs to be done; the fonts will be automatically created and installed as needed.

The downside is that it does take a certain amount of time to create a new font if it has never been used before. But once a font is created, it will exist on disk, and the next time that document is printed it will print very quickly.

It is the MakeTeXPK shell script that is responsible for making these fonts. The MakeTeXPK script supplied invokes to create the font and then copies the resultant pk file to a world-writable font cache area.

MakeTeXPK can be customized to do other things to get the font. For instance, if you are installing dvips to replace (or run alongside) an existing PostScript driver, and that driver demands gf fonts, you can easily modify MakeTeXPK to invoke gftopk to convert the gf files to pk files for dvips. This provides the same space savings listed above.

Because dvips (and thus MakeTeXPK) is run by a wide variety of users, there must be a system-wide place to put the cached font files. In order for everyone to be able to supply fonts, the directory must be world writable. If your system administrator considers this a security hole, MakeTeXPK can write to /tmp/pk or some such directory, and periodically the cached fonts can be moved to a more general system area. Note that the cache directory must exist on the pk file search path in order for MakeTeXPK to work.

11. Path Interpretation

The dvips program needs to read a wide variety of files from a large set of directories. It uses a set of paths to do this. The actual paths are listed in the next section; this section describes how the paths are interpreted.

All path variables are names of directories (path elements), separated by colons. Each path element can be either the literal name of a directory or one of the ~ forms common under UNIX. If a path element is a single tilde, it is replaced by the contents of the environment variable HOME, which is normally set to the user’s home directory. If the path element is a tilde followed by anything, the part after the tilde is interpreted as a user name, and his home directory is fetched from the system password file and used as the real path element.

Where environment variables can override paths, an additional path element form is allowed. If a path element is the empty string, it is replaced with the system defaults. Thus, to search the user’s home directory, followed by the system default paths, assuming the current shell is csh, the following command would be used:

setenv TEXINPUTS ~:

This is a path of two elements. The first is the user’s home directory. The second path element is the empty string, indicating that the system defaults should be searched.

The ‘system defaults’ as defined here means the strings set in the Makefile before compilation, rather than any defaults set in config.ps or printer-specific configuration files. This is to prevent path blowup, where more and more directories are added to the path by each level of configuration file.

12. Environment Variables

The dvips program reads a certain set of environment variables to configure its operation. The path variables are read after all configuration files are read, so they override values in the configuration files. (The TEXCONFIG variable, of course, is read before the configuration files.) The rest are read as needed.

Note that all defaults supplied here are just as supplied in the provided Makefile; they will almost certainly have been changed during installation at your particular site.

HOME (no default) This environment variable is automatically set by the shell and is used to replace any occurrences of ~ in a path.

MAKETEXPK (MakeTeXPK %n %d %b %m) This environment variable sets the command to be executed to create a missing font. A %n is replaced by the base name of the font to be created (such as cmr10); a %d is replaced by the resultant horizontal resolution of the font; a %b is replaced by the horizontal resolution at which dvips is currently generating output, a %o is replaced with the current mode, if any, or default if none is known, and any %m is replaced by a string that can use as the right hand side of an assignment to mag to create the desired font at the proper resolution. If a mode for is set in a configuration file and no %o is specified in the command, the mode is automatically appended to the command before execution. Note that these substitutions are different than the ones performed on PK paths.

DVIPSHEADERS (.:/usr/lib/tex/ps) This environment variable determines where to search for header files such as tex.pro, font files, arguments to the -h option, and such files.

PRINTER (no default) This environment variable is read to determine which default printer configuration file to read in. Note that it is the responsibility of the configuration file to send output to the proper print queue, if such functionality is desired.

TEXFONTS (/LocalLibrary/Fonts/TeXFonts/tfm:/usr/lib/tex/fonts/tfm)  This is

where tfm files are searched for. A tfm file only needs to be loaded if the font is a resident (PostScript) font or if for some reason no pk file could be found.

TEXPKS (/LocalLibrary/Fonts/TeXFonts/pk:/usr/lib/tex/fonts/pk) This environ-

ment variable is a path on which to search for pk fonts. Certain substitutions are performed if a percent sign is found anywhere in the path. See the description of the P configuration file option for more information.

TEXINPUTS (.:..:/usr/lib/tex/inputs) This environment variable is used to find PostScript figures when they are included.

TEXCONFIG (.:/usr/lib/tex/ps) This environment variable sets the directories to search for configuration files, including the system-wide one. Using this single environment variable and the appropriate configuration files, it is possible to set up the program for any environment. (The other path environment variables can thus be redundant.)

VFFONTS (.:/usr/lib/tex/fonts/vf) This environment variable sets where dvips looks for virtual fonts. A correct virtual font path is essential if PostScript fonts are to be used.

13. Other Bells And Whistles

For special effects, if any of the macros bop-hook, eop-hook, start-hook, or end-hook are defined in the PostScript userdict, they will be executed at the beginning of a page, end of a page, start of the document, and end of a document, respectively. When these macros are executed, the default PostScript coordinate system and origin is in effect. Such macros can be defined in headers added by the -h option or the header= special, and might be useful for writing, for instance, DRAFT across the entire page, or, with the aid of a shell script, dating the document. These macros are executed outside of the save/restore context of the individual pages, so it is possible for them to accumulate information, but if a document must be divided into sections because of memory constraints, such added information will be lost across section breaks.

The two arguments to bop-hook are the TEX page number and the sequence number of the page in the file; the first page gets zero, the second one, etc. The arguments to start-hook are hsize, vsize, mag, hdpi, vdpi, and the name of the dvi input file. The procedures must leave these parameters on the stack. The other hooks are not (currently) given parameters, although this may change in the future.

As an example of what can be done, the following special will write a light DRAFT across each page in the document:

    \special{!userdict begin /bop-hook{gsave 200 30 translate
    65 rotate /Times-Roman findfont 216 scalefont setfont
    0 0 moveto 0.7 setgray (DRAFT) show grestore}def end}

Note that using bop-hook or eop-hook in any way that preserves information across pages will break compliance with the Adobe document structuring conventions, so if you use any such tricks, it is recommended that you also use the -N option to turn off structured comments.

Several of the above tricks can be used nicely together, and it is not necessary that a ‘printer configuration file’ be used only to set printer defaults. For instance, you might have a file config.dated that contains just the two lines

    E echo /bop-hook \{save /Times-Roman findfont 7 scalefont setfont \
                      72 756 moveto \(`date`\) show restore\} def >.date
    h .date

(with no newline following setfont); then the command-line option -Pdated to dvips will print current date and time on the top of each page of output. Note that multiple -P options can be used.

14. Installation

If dvips has not already been installed on your system, the following steps are all that are needed.

First update the Makefile—in particular, the paths. Everything concerning dvips can be adjusted in the Makefile. Make sure you set key parameters such as the default resolution, and make sure that the path given for packed pixel files is correct.

Next, check the file name definitions in MakeTeXPK. If you don’t have installed, you cannot use MakeTeXPK to automatically generate the fonts; you can, however, modify it to generate pk fonts from gf fonts if you don’t have a full set of pk fonts but do have a set of gf fonts. If you don’t have that, you should probably not install MakeTeXPK at all; this will disable automatic font generation.

Now, check the configuration parameters in config.ps. You should also update the default resolution here. This file is the system-wide configuration file that will be automatically installed. If you are unsure how much memory your PostScript printer has, print the following file:

    %!  Hey, we're PostScript
    /Times-Roman findfont 30 scalefont setfont 144 432 moveto
    vmstatus exch sub 40 string cvs show pop showpage

Note that the number returned by this file is the total memory free; it is often a good idea to tell dvips that the printer has somewhat less memory. This is because many programs download permanent macros that can reduce the memory in the printer. In general, a memory size of about 300000 is plenty, since dvips can automatically split a document if required. It is unfortunate that PostScript printers with much less virtual memory still exist. Some systems or printers can dynamically increase the memory available to a PostScript interpreter; for these systems, a value of one million works well.

Next, run make. Everything should compile smoothly. You may need to adjust the compiler options in the Makefile if something goes amiss.

Once everything is compiled, run make install. After this is done, you may want to create a configuration file for each PostScript printer at your site.

If the font caching is considered a security hole, make the ‘cache’ directory be something like /tmp/pks, and cron a job to move the good pk files into the real directory. Or simply disable this feature by not installing MakeTeXPK.

Don’t forget to install the new vf files and tfm files. Note that the tfm files distributed with earlier (pre-5.471) versions of dvips, and all versions of other PostScript drivers, are different.

A test program called test.tex is provided, so you can easily check that the most important parts of dvips have been installed properly.

15. Diagnosing Problems

You’ve gone through all the trouble of installing dvips, carefully read all the instructions in this manual, and still can’t get something to work. This is all too common, and is usually caused by some broken PostScript application out there. The following sections provide some helpful hints if you find yourself in such a situation.

In all cases, you should attempt to find the smallest file that causes the problem. This will not only make debugging easier, it will also reduce the number of possible interactions among different parts of the system.

15.1 Debug Options

The -d flag to dvips is very useful for helping to track down certain errors. The parameter to this flag is an integer that tells what errors are currently being tracked. To track a certain class of debug messages, simply provide the appropriate number given below; if you wish to track multiple classes, sum the numbers of the classes you wish to track. The classes are:

1 specials
2 paths
4 fonts
8 pages
16 headers
32 font compression
64 files
128 memory

15.2 No Output At All

If you are not getting any output at all, even from the simplest one-character file (for instance, \ \bye), then something is very wrong. Practically any file sent to a PostScript laser printer should generate some output, at the very least a page detailing what error occurred, if any. Talk to your system administrator about downloading a PostScript error handler. (Adobe distributes a good one called ehandler.ps.)

It is possible, especially if you are using non-Adobe PostScript, that your PostScript interpreter is broken. Even then it should generate an error message. I’ve tried to work around as many bugs as possible in common non-Adobe PostScript interpreters, but I’m sure I’ve missed a few.

If dvips gives any strange error messages, or compilation on your machine generated a lot of warnings, perhaps the dvips program itself is broken. Carefully check the types in dvips.h and the declarations in the Makefile, and try using the debug options to determine where the error occurred.

It is possible your spooler is broken and is misinterpreting the structured comments. Try the -N flag to turn off structured comments and see what happens.

15.3 Output Too Small or Inverted

If some documents come out inverted or too small, your spooler is not supplying an end of job indicator at the end of each file. (This happens a lot on small machines that don’t have spoolers.) You can force dvips to do this with the -F flag, but note that this generates files with a binary character (control-D) in them. You can also try using the -s flag to enclose the entire job in a save/restore pair.

15.4 Error Messages From Printer

If your printer returns error messages, the error message gives very good information on what might be going wrong. One of the most common error messages is bop undefined. This is caused by old versions of Transcript and other spoolers that do not properly parse the setup section of the PostScript. To fix this, turn off structured comments with the -N option, but make sure you get your spooling software updated. You might also try turning off comments on included files with the -K option; many spoolers cannot deal with nested documents.

Another error message is VM exhausted. (Some printers indicate this error by locking up; others quietly reset.) This is caused by telling dvips that the printer has more memory than it actually does, and then printing a complicated document. To fix this, try lowering the parameter to m in the configuration file; use the debug option to make sure you adjust the correct file.

Other errors may indicate that the graphics you are trying to include don’t nest properly in other PostScript documents, or any of a number of other possibilities. Try the output on a QMS PS-810 or other Adobe PostScript printer; it might be a problem with the printer itself.

15.5 400 DPI Is Used Instead Of 300 DPI

This common error is caused by not editing the config.ps file to reflect the correct resolution for your site. You can use the debug flags (-d64) to see what files are actually being read.

15.6 Long Documents Fail To Print

This is usually caused by incorrectly specifying the amount of memory the printer has in config.ps; see the description above.

15.7 Including Graphics Fails

The reasons why graphics inclusions fail are too numerous to mention. The most common problem is an incorrect bounding box; read the section on bounding boxes and check your PostScript file. Complain very loudly to whoever wrote the software that generated the file if the bounding box is indeed incorrect.

Another possible problem is that the figure you are trying to include does not nest properly; there are certain rules PostScript applications should follow when generating files to be included. The dvips program includes work-arounds for such errors in Adobe Illustrator and other programs, but there are certainly applications that haven’t been tested.

One possible thing to try is the -K flag, to strip the comments from an included figure. This might be necessary if the PostScript spooling software does not read the structuring comments correctly. Use of this flag will break graphics from some applications, though, since some applications read the PostScript file from the input stream looking for a particular comment.

Any application which generates graphics output containing raw binary (not hex) will probably fail with dvips.

15.8 Can’t Find Font Files

If dvips complains that it cannot find certain font files, it is possible that the paths haven’t been set up correctly for your system. Use the debug flags to determine precisely what fonts are being looked for and make sure these match where the fonts are located on your system.

15.9 Can’t Generate Fonts

This happens a lot if either MakeTeXPK hasn’t been properly edited and installed, or if the local installation of isn’t correct. If MakeTeXPK isn’t properly edited or isn’t installed, an error such as MakeTeXPK not found will be printed on the console. The fix is to talk to the person who installed dvips and have them fix this.

If isn’t found when MakeTeXPK is running, make sure it is installed on your system. The person who installed TEX should be able to install easily.

If runs but generates fonts that are too large (and prints out the name of each character as well as just a character number), then your base file probably hasn’t been made properly. To make a proper plain.base, assuming the local mode definitions are contained in local.mf (on the NeXT, next.mf; on the Amiga, amiga.mf), type the following command (assuming csh under UNIX):

    localhost> inimf "plain; input local; dump"

Now, copy the plain.base file from the current directory to where the base files are stored on your system.

Note that a preloaded cmbase.base should never be used when creating fonts, and a program such as cmmf should never exist on the system. The macros defined in cmbase will break fonts that do not use cmbase; such fonts include the LaTEX fonts. Loading the cmbase macros when they are needed is done automatically and takes less than a second— an insignificant fraction of the total run time of for a font, especially when the possibility of generating incorrect fonts is taken into account. If you create the LaTEX font circle10, for instance, with the cmbase macros loaded, the characters will have incorrect widths.

16. Using Color with dvips

This new feature of dvips is somewhat experimental so your experiences and comments are welcome. Initially added by Jim Hafner, IBM Research, hafner@almaden.ibm.com, the color support has gone through many changes by Tomas Rokicki. Besides the changes to the source code itself, there are additional TEX macro files: colordvi.tex and blackdvi.tex. There are also .sty versions of these files that can be used with LaTEX and other similar macro packages. This feature adds one-pass multi-color printing of TEX documents on any color PostScript device.

In this section we describe the use of color from the document preparer’s point of view and then add some simple instructions on installation for the system administrator.

16.1 The Macro Files

All the color macro commands are defined in colordvi.tex (or colordvi.sty). To access these macros simply add to the top of your TEX file the command

    \input colordvi

or, if your document uses style files like LaTEX, add the colordvi style option as in

    \documentstyle[12pt,colordvi]{article}

There are basically two kinds of color macros, ones for local color changes to, say, a few words or even one symbol and one for global color changes. Note that all the color names use a mixed case scheme. There are 68 predefined colors, with names taken primarily from the Crayola crayon box of 64 colors, and one pair of macros for the user to set his own color pattern. More on this extra feature later. You can browse the file colordvi.tex for a list of the predefined colors. The comments in this file also show a rough correspondence between the crayon names and PANTONEs.

A local color command is in the form

    \ColorName{this will print in color}

Here ColorName is the name of a predefined color. As this example shows, this type of command takes one argument which is the text that is to print in the selected color. This can be used for nested color changes since it restores the original color state when it completes. For example, suppose you were writing in green and want to switch temporarily to red, then blue, back to red and restore green. Here is one way that you can do this:

    This text is green but here we are \Red{switching to red,
    \Blue{nesting blue}, recovering the red} and back to
    original green.

In principle there is no limit to the nesting level, but it is not advisable to nest too deep lest you lose track of the color history.

The global color command has the form

    \textColorName

This macro takes no arguments and immediately changes the default color from that point on to the specified color. This of course can be overriden globally by another such command or locally by local color commands. For example, expanding on the example above, we might have

    \textGreen
    This text is green but here we are \Red{switching to red,
    \Blue{nesting blue}, recovering the red} and back to
    original green.
    \textCyan
    The text from here on will be cyan unless
    \Yellow{locally changed to yellow}.  Now we are back to cyan.

The color commands will even work in math mode and across math mode boundaries. This means that if you have a color before going into math mode, the mathematics will be set in that color as well. More importantly however, in alignment environments like \halign, tabular or eqnarray, local color commands cannot extend beyond the alignment characters.

Because local color commands respect only some environment and delimiter changes besides their own, care must be taken in setting their scope. It is best not to have then stretch too far.

At the present time there are no macros for color environments in LaTEX which might have a larger range. This is primarily to keep the TEX and LaTEX use compatible.

16.2 User Definable Colors

There are two ways for the user to specify colors not already defined. For local changes, there is the command \Color which takes two arguments. The first argument is a quadruple of numbers between zero and one and specifies the intensity of cyan, magenta, yellow and black (CMYK) in that order. The second argument is the text that should appear in the given color. For example, suppose you want the words “this color is pretty” to appear in a color which is 50% cyan, 85% magenta, 40% yellow and 20% black. You would use the command

    \Color{.5 .85 .4 .2}{this color is pretty}

For global color changes, there is a command \textColor which takes one argument, the CMYK quadruple of relative color intensities. For example, if you want the default color to be as above, then the command

    \textColor{.5 .85 .4 .2}
    The text from now on will be this pretty color

will do the trick.

Making a global color change in the midst of nested local colors is highly discouraged. Consequently, dvips will give you warning message and do its best to recover by discarding the current color history.

16.3 Subtleties in Using Color

These color macros are defined by use of specialized \special keywords. As such, they are put in the .dvi file only as explicit message strings to the driver. The (unpleasant) result is that certain unprotected regions of the text can have unwanted color side effects. For example, if a color region is split by TEX across a page boundary, then the footers of the current page (e.g., the page number) and the headers of the next page can inherit that color. To avoid this effect globally, users should make sure that these special regions of the text are defined with their own local color commands. For example in TEX, to protect the header and footer, use

    \headline{\Black{My Header}}
    \footline{\Black{\hss\tenrm\folio\hss}}

This warning also applies to figures and other insertions, so be careful!

Of course, in LaTEX, this is much more difficult to do because of the complexity of the macros that control these regions. This is unfortunate, but is somehow inevitable because TEX and LaTEX were not written with color in mind.

Even when writing your own macros, much care must be taken. The color macros that ‘colorize’ a portion of the text work by prefixing the text with a special command to turn the color on and postfixing it with a different special command to restore the original color. It is often useful to insure that TEX is in horizontal mode before the first special command is issued; this can be done by prefixing the color command with \leavevmode.

16.4 Printing in Black/White, after Colorizing

If you have a TEX or LaTEX document written with color macros and you want to print it in black and white there are two options. On all (good) PostScript devices, printing a color file will print in corresponding grey-levels. This is useful since in this way you can get a rough idea of where the colors are changing without using expensive color printing devices. The second option is to replace the call to input colordvi.tex with blackdvi.tex (and similarly for the .sty files). So in the above example, replacing the word colordvi with blackdvi suffices. This file defines the color macros as no-ops, and so will produce normal black/white printing. By this simple mechanism, the user can switch to all black/white printing without having to ferret out the color commands. Also, some device drivers, particularly non-PostScript ones like screen previewers, will simply ignore the color commands and so print in normal black/white. Hopefully, in the future screen previewers for color displays will be compatible with some form of color support.

16.5 Configuring dvips for Color Devices

To configure dvips for a particular color device you need to fine tune the color parameters to match your device’s color rendition. To do this, you will need a PANTONE chart for your device. The header file color.lpro shows a (rough) correspondence between the Crayola crayon names and the PANTONE numbers and also defines default CMYK values for each of the colors. Note that these colors must be defined in CMYK terms and not RGB as dvips outputs PostScript color commands in CMYK. This header file also defines (if they are not known to the interpreter) the PostScript commands setcmykcolor and currentcmykcolor in terms of a RGB equivalent so if your device only understands RGB, there should be no problem.

The parameters set in this file were determined by comparing the PANTONE chart of a Tektronics PHASER printer with the actual Crayola Crayons. Because these were defined for a particular device, the actual color rendition on your device may be very different. There are two ways to adjust this. One is to use the PANTONE chart for your device to rewrite color.lpro prior to compilation and installation. A better alternative, which supports multiple devices, is to add a header file option in the configuration file for each device that defines, in userdict, the color parameters for those colors that need redefining.

For example, if you need to change the parameters defining Goldenrod (approximately PANTONE 109) for your device mycolordev, do the following. In the PANTONE chart for your device, find the CMYK values for PANTONE 109. Let’s say they are { 0 0.10 0.75 0.03 }. Then create a header file named mycolordev.pro with the commands

    userdict begin
    /Goldenrod { 0 0.10 0.75 0.03 setcmykcolor} bind def

Finally, in config.mycolordev add the line

    h mycolordev.pro

This will then define Goldenrod in your device’s CMYK values in userdict which is checked before defining it in TeXdict by color.pro.

This mechanism, together with additions to colordvi.tex and blackdvi.tex (and the .sty files), can also be used to predefine other colors for your users.

16.6 Protecting Regions From Spurious Colors

Because color is defined via TEX’s \special command, it cannot be sensitive to the output routine or certain regions of the page like the header and footer. Consequently, these regions need to be protected from spurious color changes (particularly when local colors spread across page breaks).

Users need to be aware of the possibility of certain regions getting unwanted or unpredicted colors. Headers and footers are most worrisome so style designers who want to use color should keep this in mind.

One particular region of text that gets spurious color effects is labels in list environments. For instance, because of the way list items are defined in standard LaTEX, the bullet for items that start with a different color also gets drawn in that color.

To give the user a simple mechanism to solve this problem (and other unforeseen effects of this type) one other special macro is automatically defined. This macro is called \globalColor. It is actually a local color macro and so takes a single argument. But the color effect it produces is always the same as that set by the last \textColor or \textColorName command. In effect, when a \textColorName command is called, \globalColor gets a new definition equivalent to the local \ColorName macro. For example, when the default is

black, \globalColor=\Black and when \textGreen appears, \globalColor=\Green. This

special macro can then be used to protect sensitive regions of the text.

For example, in LaTEX files, one might make sure that the header and footers have \globalColor wrapping their contents. In this way, they will inherit the current active root (unnested) color state.

16.7 Color Support Details

To support color, dvips recognizes a certain set of specials. These specials all start with the keyword color or the keyword background.

We will describe background first, since it is the simplest. The background keyword must be followed by a color specification. That color specification is used as a fill color for the background. The last background special on a page is the one that gets issued, and it gets issued at the very beginning of the page, before any text or specials are sent. (This is possible because the prescan phase of dvips notices all of the color specials so that the appropriate information can be written out during the second phase.)

Ahh, but what is a color specification? It is one of three things. First, it might be a PostScript procedure as defined in a PostScript header file. The color.pro file defines 64 of these, including Maroon. This PostScript procedure must set the current color to be some value; in this case, Maroon is defined as 0 0.87 0.68 0.32 setcmykcolor.

The second possibility is the name of a color model (initially, one of rgb, hsb, cmyk, or gray) followed by the appropriate number of parameters. When dvips encounters such a macro, it sends out the parameters first, followed by the string created by prefixing TeXcolor to the color model. Thus, the color specification rgb 0.3 0.4 0.5 would generate the PostScript code 0.3 0.4 0.5 TeXrgbcolor. Note that the case of zero arguments is disallowed, as that is handled by the single keyword case above (where no changes to the name are made before it is sent to the PostScript file.)

The third and final type of color specification is a double quote followed by any sequence of PostScript. The double quote is stripped from the output. For instance, the color specification "AggiePattern setpattern will set the ‘color’ to the Aggie logo pattern (assuming such exists.)

So those are the three types of color specifications. The same type of specifications are used by both the background special and the color special. The color special itself has three forms. The first is just color followed by a color specification. In this case, the current global color is set to that color; the color stack must be empty when such a command is executed.

The second form is color push followed by a color specification. This saves the current color on the color stack and sets the color to be that given by the color specification. This is the most common way to set a color.

The final version of the color special is just color pop, with no color specification; this says to pop the color last pushed on the color stack from the color stack and set the current color to be that color.

The dvips program correctly handles these color specials across pages, even when the pages are repeated or reversed.

These color specials can be used for things such as patterns or screens as well as simple colors. However, note that in the PostScript, only one ‘color specification’ can be active at a time. For instance, at the beginning of a page, only the bottommost entry on the color stack is sent; also, when a color is ‘popped’, all that is done is that the color specification from the previous stack entry is sent. No gsave or grestore is used. This means that you cannot easily mix usage of the color specials for screens and colors, just one or the other. This may be addressed in the future by adding support for different ‘categories’ of color-like state.